# Benchmarking direct and indirect dipolar spin-exchange interactions between two Rydberg atoms

### v2

@@ I Experimental setup
Before the Rydberg sequence, we use Raman sideband cooling [33, 34] to reduce the thermal fluctuations of atomic motion inside the tweezers, achieving typical average occupation numbers for the motional states of $\bar{n}_{\perp}\sim 0.5$ radially, and $\bar{n}_{\parallel}\sim 10$ axially. The standard deviation of the position along a direction $i$ depends on the tweezer trapping frequency $\omega_{i}$ as $\sigma_{i}=\sqrt{\hbar(\bar{n}_{i}+1/2)/m\omega_{i}}$, and the standard deviation of the velocity distribution as $\sigma_{v_{i}}=\omega_{i}\sigma_{i}$. The total position uncertainty after a duration $t$ of free flight is thus $\sigma_{i}(t)=\sqrt{\sigma_{i}^{2}+(\sigma_{v_{i}}t)^{2}}$, as pictured in Fig. 1(a). Before the Rydberg sequence, we adiabatically reduce the trapping depth by a factor $\zeta\in[0.001,1]$, and thus the trapping frequencies by a factor $\sqrt{\zeta}$ (starting from $\omega_{\perp}\sim 2\pi\times 100$ kHz and $\omega_{\parallel}\sim 2\pi\times 20$ kHz). This corresponds to a positional disorder of $\sigma_{\perp}\sim[30,200]$ nm, $\sigma_{\parallel}\sim[250,1400]$ nm, $\sigma_{v_{\perp}}\sim[4,20]$ nm///µs and $\sigma_{v_{\parallel}}\sim[6,30]$ nm///µs, leading in turn to fluctuations of the interatomic distance. To first order in $\sigma_{\perp}$ and $\sigma_{\parallel}$, the standard deviation of distances is given by $\Delta r(t)=\sqrt{2}\sigma_{\perp}(t)$, which is a few percents of the average distance $r$.
@@ V Outlook
In this work, we have illustrated, in the simplest setting of just two atoms at a controlled distance, how various types of spin-exchange couplings can be implemented by encoding two spin states in Rydberg levels. Our careful benchmarking of the observed coherent interactions highlights the effects of positional disorder arising from the motion of the atoms, and that limits coherence to 10-20 interaction times. A possible way to mitigate these detrimental effects is to increase the interparticle distance $r$ and use Rydberg states with a larger principal quantum number, such that the relative fluctuations of interaction energy are reduced as $\Delta J/|J|\propto\Delta r/r$. Another solution consists in trapping the Rydberg states using ponderomotive potentials [47]; however recent studies [48] suggest that one would need to operate very far in the Lamb-Dicke regime to make the effect of zero-point fluctuations entirely negligible.
In the experiments reported here, the motion of the atoms leads to dephasing of the spin-exchange oscillations but, since we restrict ourselves to relatively short times, the dipolar interaction does not influence the motion of the atoms (as discussed in App. A). However, going one step further, one can also take into account the mechanical effect of van der Waals or dipolar interactions on the motion of the atoms. This gives rise to a spin-motion coupling, observed for instance in [7, 49], that yields entanglement between internal and external degrees of freedom. From the perspective of quantum simulation of spin models, this is yet another decoherence channel; however, this spin-motion coupling may also open interesting perspectives. For example, considering that the relative atomic motion is described by a one dimensional harmonic oscillator with annihilation and creation operators $\hat{a},\hat{a}^{\dagger}$: $\hat{r}=r_{0}+\Delta x(\hat{a}+\hat{a}^{\dagger})$, the Taylor expansion of Eq. (1) contains a term of the form $(|J|\Delta x/r)(\hat{a}+\hat{a}^{\dagger})(\hat{\sigma}_{1}^{+}\hat{\sigma}_{2}^{-}+\hat{\sigma}_{1}^{-}\hat{\sigma}_{2}^{+})$ (with $\hat{\sigma}^{\pm}$ the pseudo-spin operators), which could be tuned. Spin-motion coupling has also be suggested to implement GKP codes [50] for quantum error correction.
@@ Appendix B A simple analytic model for the damping of a spin exchange oscillation
The envelope shows periodic revivals due to the oscillations of the atomic positions in the traps, which cause the probability of each classical trajectory [Eq. (16)] to rephase. Those revivals were also derived in Ref. [48] using a full quantum treatment of the atomic motion (see Appendix C). The modulation of the envelope in Eq. (21) can be suppressed by increasing the trapping frequency $\omega_{\perp}$ so to reach the regime $\omega_{\perp}\gg\Delta J$. Using the dependence of $\Delta J$ on $\omega_{\perp}$, the condition to suppress the modulation of the envelope becomes $\omega_{\perp}\gg\omega_{\perp}^{\mathrm{crit}}$ with a critical trapping frequency
$\omega_{\perp}^{\mathrm{crit}}=\left[\frac{\alpha J_{0}}{r_{0}}\sqrt{\frac{\hbar(2\bar{n}_{\perp}+1)}{m}}\right]^{2/3}.$ (22)
For the parameters of Fig. 2(c), we find $\omega_{\perp}^{\mathrm{crit}}=2\pi\times 11$ kHz in the ground state.
In Fig. 6(b, top panel), we simulate the spin exchange of two trapped atoms in their motional ground state for $\omega_{\perp}=2\pi\times 37$ kHz $>\omega_{\perp}^{\mathrm{crit}}$ and find again a good agreement with the predicted profile of Eq. (21). Now finite Rydberg lifetimes become the dominating limiting factor to the damping [Fig. 6(b, bottom panel)], leading to a significant improvement in the coherence of the oscillation compared with the case of free atoms.
To estimate by how much Rydberg trapping can improve the damping, we repeat the simulations for various initial trapping frequencies $\omega_{\perp}$ and extract the quality factor of each resulting spin-exchange oscillation. Due to the revivals, fitting $P_{\uparrow\downarrow}(t)$ by a Gaussian envelope is not reliable, so we define the damping time $\tau$ as the $1/\sqrt{e}$ decay time for the contrast of $P_{\uparrow\downarrow}(t)$. The results are shown in Fig. 6(c) as orange lines: the quality factor reaches $\sim 100$ for trapped Rydberg atoms and saturates starting from $\omega_{\perp}\sim 2\pi\times 37$ kHz due to the finite lifetimes.

### v3

@@ I Experimental setup
Before the Rydberg sequence, we use Raman sideband cooling [33, 34] to reduce the thermal fluctuations of atomic motion inside the tweezers, achieving typical average occupation numbers for the motional states of $\bar{n}_{\perp}\sim 0.5$ radially, and $\bar{n}_{\parallel}\sim 10$ axially. The corresponding temperatures are $T_{\perp}\sim 4$ µK and $T_{\parallel}\sim 10$ µK. The standard deviation of the position along a direction $i$ depends on the tweezer trapping frequency $\omega_{i}$ as $\sigma_{i}=\sqrt{\hbar(\bar{n}_{i}+1/2)/m\omega_{i}}$, and the standard deviation of the velocity distribution as $\sigma_{v_{i}}=\omega_{i}\sigma_{i}$. The total position uncertainty after a duration $t$ of free flight is thus $\sigma_{i}(t)=\sqrt{\sigma_{i}^{2}+(\sigma_{v_{i}}t)^{2}}$, as pictured in Fig. 1(a). Before the Rydberg sequence, we adiabatically reduce the trapping depth by a factor $\zeta\in[0.001,1]$, and thus the trapping frequencies by a factor $\sqrt{\zeta}$ (starting from $\omega_{\perp}\sim 2\pi\times 100$ kHz and $\omega_{\parallel}\sim 2\pi\times 20$ kHz). This corresponds to a positional disorder of $\sigma_{\perp}\sim[30,200]$ nm, $\sigma_{\parallel}\sim[250,1400]$ nm, $\sigma_{v_{\perp}}\sim[4,20]$ nm///µs and $\sigma_{v_{\parallel}}\sim[6,30]$ nm///µs, leading in turn to fluctuations of the interatomic distance. To first order in $\sigma_{\perp}$ and $\sigma_{\parallel}$, the standard deviation of distances is given by $\Delta r(t)=\sqrt{2}\sigma_{\perp}(t)$, which is a few percents of the average distance $r$.
@@ V Outlook
In this work, we have illustrated, in the simplest setting of just two atoms at a controlled distance, how various types of spin-exchange couplings can be implemented by encoding two spin states in Rydberg levels. Our careful benchmarking of the observed coherent interactions highlights the effects of positional disorder arising from the motion of the atoms, and that limits coherence to 10-20 interaction times. A possible way to mitigate these detrimental effects is to increase the interparticle distance $r$ and use Rydberg states with a larger principal quantum number, such that the relative fluctuations of interaction energy are reduced as $\Delta J/|J|\propto\Delta r/r$. Another solution consists in trapping the Rydberg states using ponderomotive potentials [47], which can eliminate the effect of positional disorder when the trapping frequency $\omega_{\perp}$ dominates over the fluctuations of interaction energy $\Delta J$ [48] (see also App. B).
In the experiments reported here, the motion of the atoms leads to dephasing of the spin-exchange oscillations but, since we restrict ourselves to relatively short times, the dipolar interaction does not influence the motion of the atoms (as discussed in App. A). However, going one step further, one can also take into account the mechanical effect of van der Waals or dipolar interactions on the motion of the atoms. This gives rise to a spin-motion coupling, observed for instance in [7, 49], that yields entanglement between internal and external degrees of freedom. From the perspective of quantum simulation of spin models, this is yet another decoherence channel; however, this spin-motion coupling may also open interesting perspectives, for example as a resource for generating entangled states [50, 48, 51, 52]. It has also been suggested to implement GKP codes [53] for quantum error correction.
@@ Appendix B A simple analytic model for the damping of a spin exchange oscillation
The envelope shows periodic revivals due to the oscillations of the atomic positions in the traps, which cause the probability of each classical trajectory [Eq. (16)] to rephase. Those revivals were also derived in Ref. [48] using a full quantum treatment of the atomic motion (see Appendix C). The modulation of the envelope in Eq. (21) can be suppressed by increasing the trapping frequency $\omega_{\perp}$ so to reach the regime $\omega_{\perp}\gg\Delta J$. Using the dependence of $\Delta J$ on $\omega_{\perp}$, this condition becomes $\omega_{\perp}\gg\omega_{\perp}^{\mathrm{opt}}/3$.
In Fig. 6(b, top panel), we simulate the spin exchange of two trapped atoms in their motional ground state for $\omega_{\perp}\simeq\omega_{\perp}^{\mathrm{opt}}$ and find again a good agreement with the predicted profile of Eq. (21). Now finite Rydberg lifetimes become the dominating limiting factor to the damping [Fig. 6(b, bottom panel)], leading to a significant improvement in the coherence of the oscillation compared with the case of free atoms.
To estimate by how much Rydberg trapping can improve the damping, we repeat the simulations for various initial trapping frequencies $\omega_{\perp}$ and extract the quality factor of each resulting spin-exchange oscillation. Due to the revivals, fitting $P_{\uparrow\downarrow}(t)$ by a Gaussian envelope is not reliable, so we define the damping time $\tau$ as the $1/\sqrt{e}$ decay time for the contrast of $P_{\uparrow\downarrow}(t)$. The results are shown in Fig. 6(c) as orange lines: the quality factor reaches $\sim 100$ for trapped Rydberg atoms and saturates starting from $\omega_{\perp}\sim\omega_{\perp}^{\mathrm{opt}}$ due to the finite lifetimes.